OpenStudy (unknownrandom):

Can somebody show me how to work this out? Find all of the zeros f(x) = 48x^4 - 52x^3 + 13x - 3

OpenStudy (mathmale):

Dear UnknownRandom: Do you happen to know and understand how to use synthetic division?

OpenStudy (unknownrandom):

Yes

OpenStudy (anonymous):

(52x(3)+13x)−3 =52x(3)+13x+(−3) =169x+(−3) =169x−3

OpenStudy (anonymous):

Please give a medal and fan if this was helpful! :D

OpenStudy (mathmale):

Look at the first and last coefficients of the given polynomial: 48 and -3. Multiply those together and then look for pairs of factors. -3(48)=-144. What are factors of -144? This method depends somewhat on trial and error, but after you've identified some factors of -144 you could try them, one by one, as divisors in synth. div. Nikkeyboo: I'd think medals and "fanning" would mean more if given spontaneously and not requested.

OpenStudy (mertsj):

Easiest way is to use synthetic division and put 1 or -1 on the shelf.

OpenStudy (mathmale):

I think so too. But what do you mean by "put 1 or -1 on the shelf?"

OpenStudy (unknownrandom):

@mathmale That makes a lot of since. @Mertsj Alright I'll try that.

OpenStudy (mertsj):

.5 | 48 -52 0 13 -3 24 -14 -7 3 48 -28 -14 6

OpenStudy (mertsj):

Actually, it has to be 1/2 or -1/2. Sorry about that

OpenStudy (mertsj):

So if 1/2 is a 0 then 2x-1 is a factor.

OpenStudy (mathmale):

Nice work, Mertsj! Please explain 2 things to UR: how you arrived at that .5, and at that -.5, and how you knew from your example of synth. div. that .5 is a root. I'm interested in knowing your reasoning.

OpenStudy (mathmale):

Note, PP and Mertsj, that 1/2 has numerator and denominator; that 1 in the numerator is a factor of 3, whereas that 2 in the denominator is a factor of 48 (but not of -3). Mertsj: Can you build on this explanation, please?

OpenStudy (mertsj):

The possible rational roots are the factors of the constant term which is 3 divided by the factors of the coefficient of the leading term which is 48. I tried 1 and -1 and because of the way the numbers worked out, (and lots of experience) I tried 1/2. It worked. So I tried -1/2 and it worked as well. Then I have a quadratic trinomial which can be factored by the usual methods.

OpenStudy (mathmale):

UR: I'd like to hear YOUR perspectives also! After all, it's your math problem.

OpenStudy (unknownrandom):

I see what @Mertsj is doing. I was wondering if their was a faster way to do it.

OpenStudy (mathmale):

Yes, Mertsj. To paraphrase what you've said: "Possible rational roots of the given polynomial have numerators which are factors of the last (the constant) term, and denominators which are factors of the first (the leading) term."

OpenStudy (mertsj):

Factors of 3: 1 -1 3 -3 Factors of 48: 1, -1, 2, -2, 3, -3, 4, -4, 6, -6, 8, -8, 12, -12, 16, -16, 24, -24, 48, -48, 3/6=1/2 -3/6=-1/2

OpenStudy (mathmale):

Everybody wants a faster way to do stuff. But in the long run learning how to do problems correctly and well is more important than speed any day.

OpenStudy (mathmale):

Mertsj: C O O L. Thank you.

OpenStudy (unknownrandom):

@mathmale I agree 100%. Thanks again @mathmale and @Mertsj

OpenStudy (mertsj):

yw

OpenStudy (mathmale):

Bye!